OpenStudy (anonymous):

show that 3cos^2-sin^2=1+2cos2x

hartnn (hartnn):

use, cos2x = cos^2x-sin^2x

OpenStudy (anonymous):

Solving right hand side first @hartnn

hartnn (hartnn):

yup, easier

OpenStudy (anonymous):

I will go with Left then..

hartnn (hartnn):

let asker try first

OpenStudy (anonymous):

i did use that rule but i got stuck there

hartnn (hartnn):

which step exactly? can u show your work?

hartnn (hartnn):

so that we can guide u from there

OpenStudy (anonymous):

put cos^2(x)= 1/2 (1+cos2x) and sin^2(x)= 1/2 (1 - cos2x) to get the reqd ans

OpenStudy (anonymous):

Show your work first @david warner..

OpenStudy (anonymous):

i basically did it from the LHS .. used the cos(2x) rule .. so i hhad 3(cos2x) and then (3(2cos^2x-1)

hartnn (hartnn):

thats where u did the mistake....its 3cos^2-sin^2 and not 3cos^2-3sin^2.... write cos^2(x)= 1/2 (1+cos2x) and sin^2(x)= 1/2 (1 - cos2x)

OpenStudy (anonymous):

ohh i see it now .. thanku :)

hartnn (hartnn):

welcome ^_^ ask if there are still doubts anywhere